### v1

# Weiqi Wang

Abstract
To enable Large Language Models (LLMs) to function as conscious agents with generalizable reasoning capabilities, it is crucial that they possess the reasoning ability to comprehend situational changes (transitions) in distribution triggered by environmental factors or actions from other agents. Despite its fundamental significance, this ability remains underexplored due to the complexity of modeling infinite possible changes in an event and their associated distributions, coupled with the lack of benchmark data with situational transitions. Addressing these gaps, we propose a novel formulation of reasoning with distributional changes as a three-step discriminative process, termed as MetAphysical ReaSoning. We then introduce the first-ever benchmark, [Uncaptioned image]Mars, comprising three tasks corresponding to each step. These tasks systematically assess LLMs’ capabilities in reasoning the plausibility of (i) changes in actions, (ii) states caused by changed actions, and (iii) situational transitions driven by changes in action. Extensive evaluations with 20 (L)LMs of varying sizes and methods indicate that all three tasks in this process pose significant challenges, even for state-of-the-art LLMs and LMs after fine-tuning. Further analyses reveal potential causes for the underperformance of LLMs and demonstrate that pre-training them on large-scale conceptualization taxonomies can potentially enhance their metaphysical reasoning capabilities. Our data and models are publicly accessible at https://github.com/HKUST-KnowComp/MARS.
1 Introduction
Recent advancements in LLMs have demonstrated superior performance across a variety of reasoning tasks, such as causal reasoning [41, 12, 37], commonsense reasoning [51, 10, 31], mathematical reasoning [30, 21, 1], and more. However, to truly achieve conscious processing [3], the integration of System II reasoning ability [56, 35] is essential as it enables LLMs to perform out-of-distribution generalization when encountered with unfamiliar scenarios [8]. Among several components that make up System II reasoning, a critical element of it is the ability to reason with situational changes in distribution, triggered by environmental factors and actions by themselves or other agents, when dealing with non-stationarities [6]. It serves as the core ability in planning tasks [28], which can be achieved by dynamically recombining existing concepts in the given environment or action and learning from the resultant situational changes [38, 4, 17]. For instance, in the event that “PersonX is driving a car in a sunny day,” a change in the weather from sunny to rainy could cause a different outcome, such as “PersonX becomes more cautious and drives slower.” This illustrates that a change in weather conditions can lead to a change in the driver’s behavior, which represents an environmental change that triggers situational changes within the distribution of different weathers.
Though fundamental, the exploration of this ability has been limited due to several factors. First, the scope for change within an event is vast, with numerous components capable of altering in a wide variety of ways. This results in an overwhelmingly large number of potential changes that are impossible to fully cover with existing knowledge bases. Second, reasoning with changes in distribution lacks a clear formulation due to its complexity. Unlike one-step inference reasoning tasks [54, 9], changes in action may lead to implausible events that cannot occur in reality, thus terminating the reasoning process. Lastly, there is a lack of a reliable evaluation benchmark. Existing benchmarks [62, 25] typically focus on a limited number of changes within a few scenarios, thus limiting the coverage of formed distributions. The changes in actions and states are also formulated under planning or logical tasks, which neglect transitions (consequences) caused by changes.
To address these gaps, we take a step forward by formally defining reasoning with changes in distribution as a three-step discriminative process. We start by defining seven categories of changes, each corresponding to different components within an event. To semantically cover more changes in a unified manner, we propose implementing changes by altering each component within the event using their abstractions or numerical variations. This approach creates a hierarchical distribution of various changes, with the abstracted ones offering a more generalized coverage. Inspired by [8], we formulate reasoning with changes in distribution as sequentially tasking the model to: (1) assess the plausibility of a potential change in a given event that describes an action, (2) evaluate the plausibility of an inferential state resulting from the modified action, and (3) determine the necessary change in an action to convert an implausible inferential state into a plausible one. We refer to this process as metaphysical reasoning, as it requires models to distinguish implausible actions, states, and transitions that only exist in the metaphysical realm, indicating their rare occurrence in reality [26].
We then construct the first evaluation benchmark, [Uncaptioned image]Mars, featuring 355K annotated data across three tasks corresponding to each step. It is constructed by sequentially instructing ChatGPT to extract events from Wikitext [46] and BookCorpus [79], identify mutable components within each event, generate abstractions and numerical variations for those components, create a metaphysical inference state based on the changes, and generate the necessary modifications to make the metaphysical inference plausible in reality. Large-scale human annotations are then conducted to verify the generations and construct them into an evaluation benchmark. Extensive experiments with over 20 language models demonstrate that all three tasks in this process present significant challenges, even for LMs after fine-tuning. Further analyses reveal potential reasons for such underperformance and identify possible solutions for enhancing the metaphysical reasoning abilities of language models.
In summary, the main contributions of our work are three-fold:
  • Task. We introduce the task of metaphysical reasoning, which includes three distinct subtasks, designed to assess the ability of (L)LMs to reason with changes in distribution.
  • Resource. We carefully curate a large-scale evaluation benchmark, [Uncaptioned image]Mars, to facilitate evaluations of LMs’ metaphysical reasoning abilities. The dataset is released.
  • Evaluation. We experiment with several (L)LMs to demonstrate the difficulty of our proposed tasks and conduct analysis to identify the reasons behind their underperformance.
2 Backgrounds and Related Works
Reasoning about Changes in Distribution. Enabling LMs to understand distributional changes due to localized causal interventions, particularly in semantic spaces, has long been a crucial objective in the pursuit of conscious machine intelligence [7, 8]. Previous works have mainly explored this within the context of discriminating changes between actions and states with methods such as commonsense knowledge injection [58], event calculus [5], and fuzzy reasoning [77]. Other studies aim to benchmark this reasoning process through logical reasoning tasks [25] and planning tasks [62, 71]. However, these studies only cover changes in limited formats and scenarios and also overlook the significance of representing them as a distribution in relation to different variables in actions. This restricts the out-of-distribution generalizability of the resulting LMs when facing unfamiliar scenarios. Moreover, previous evaluations do not cover transitions caused by changes, making subsequent evaluations incomplete. We address these issues by proposing to use the abstraction or numerical variations of components as changes to form generalizable distributions. We also design a task in Mars to evaluate LMs’ proficiency in understanding how changes motivate situational transitions.
Benchmarking LLMs. The advent of LLMs [48, 49, 61, 60, 53] has sparked various studies in investigating LLM’s potential in a variety of tasks, including text generation [45, 14, 13], temporal reasoning [57, 76], causal reasoning [12, 15, 34], commonsense reasoning [31, 10], and more [51]. These studies have significantly contributed to our understanding of LLMs by evaluating their performance across diverse tasks, using different scales of parameters and prompting methods [50]. However, there is an absence of a comprehensive benchmark for assessing the ability of (L)LMs to reason with changes in distribution. This inspires us to formally define it and introduce the first benchmark that evaluates such reasoning capabilities of (L)LMs.
3 Definitions of Changes in Event and Metaphysical Reasoning
Figure 1: Examples of changes in event in our formulation. Events may become metaphysical as components are abstracted into high-level concepts, while some remain plausible in reality.
Modeling changes within an event is inherently complex due to the infinite combinations that can occur. For simplicity, we only consider events that represent an action and changes between their inferential states. Given an event $e$, we first define seven types of changes that could transpire within $e$. These changes are represented as components of the event, including its subject $s$, verb $v$, object $o$, temporal quantifier $t$, spatial quantifier $l$, numerical properties $n$, and sub-events $se$. The original event is denoted as a function of these seven components, $e=f(s,v,o,t,l,n,se)$. A change in the event can be represented by altering one of its components, for instance, $e^{\prime}=f(s^{\prime},v,o,t,l,n,se)$ if the change impacts the subject $s^{\prime}$.
To effectively model the distribution of changes across different types of components, we leverage two types of hierarchical formulations. Specifically, for $s,v,o,se$, we define changes in these components as conceptualizing their original instance into three concepts with progressively increased abstractedness [22, 59]. For $t,l,n$, we define their changes as modifications from their original values to three distinct numerical or spatial values with progressively increased units. This brings a hierarchical structure to changes of a certain component, forming a distribution that gradually covers more possible changes. Abstracted conponents, as high-level concepts, can semantically represent a broader range of combinations for altering an event. Some running examples of how changes impact an action are shown in Figure 1.
We then propose a three-step discriminative process, which we term as Metaphysical Reasoning, to formulate reason with changes in distribution. The three steps, as shown in Figure 2, are:
(1) Metaphysical Event Discrimination: The first step answers the question, “Will the change happen in reality?” It aims to determine the plausibility of a change based on a given event, as alterations in components may lead to implausible events that defy reality. We refer to such an event, which rarely occurs in reality due to these changes, as a metaphysical event. The goal of the first task is to discriminate whether the modified event $e^{\prime}$, conditioned on the original event e with a single altered component $c\in(s,v,o,t,l,n,se)$, is metaphysical or not by making a binary prediction.
Figure 2: Three steps in metaphysical reasoning. Our motivation behind this is that, by conquering all steps sequentially, a conscious agent could answer: (1) Will the change occur in reality? (2) What will the change cause? (3) What change can make a metaphysical (desired) inference plausible?
(2) Metaphysical Inference Discrimination: Considering that distributional changes occur in non-stationary environments, a conscious agent should be able to predict the potential outcomes of the modified event for future reasoning scenarios. Therefore, the second step aims to answer the question, “What will the altered event result in?” Similarly, we term the inferences of an event that rarely occurs in reality as metaphysical inference. The objective of the second task is to determine whether an inferential state $i$, triggered by the altered event $e^{\prime}$, is metaphysical or not by predicting a binary answer. Note that $e^{\prime}$ could be either metaphysical or not, as inferences in both cases can be evaluated.
(3) Metaphysical Transition Reasoning: Finally, with some inferences remain metaphysical, a conscious agent should be able to plan what change is necessary to make such inference plausible in reality. This completes the reasoning chain by covering the feasibility, consequence, and motivation of distributional changes. Thus, the last task answers the question, “What change is needed to make a metaphysical inference plausible?” We refer to this as metaphysical transition reasoning and set the objective as to determine whether another change, denoted as $c^{\prime}$, can make a metaphysical inference $i$ plausible in relation to a changed event $e^{\prime}$ by making a binary prediction regarding $c^{\prime}$.
4 [Uncaptioned image]Mars Benchmark Curation Pipeline
We then introduce our sequential pipeline for curating the [Uncaptioned image]Mars benchmark. An overview of our curation pipeline is shown in Figure 3. To guarantee a comprehensive coverage of events across various domains and topics in Mars, we source original text from two publicly available large corpora: Wikitext [46] and BookCorpus [79]. We filter out noisy text that includes hashtags, hyperlinks, and the like, and segment long text into sentences, each comprising no more than 200 tokens, to facilitate future processing.
4.1 Text Decomposition and Component Extraction
We first perform text decomposition [74, 32] to break down lengthy text into semantically complete short events, which are then used for fine-grained component extraction. To enable large-scale processing, we use ChatGPT [48], a powerful LLM with strong text understanding abilities, as the core processor for all stages. For each stage, we guide it with a few-shot prompt [69, 11] by creating task-specific explanations and exemplars (more details about prompts are in Appendix C):
<TASK-PROMPT>
<INPUT1><OUTPUT(1,1)>  …  <OUTPUT${}_{(1,N_{1})}$>
<INPUT2><OUTPUT(2,1)>  …  <OUTPUT${}_{(2,N_{2})}$>
<INPUT10><OUTPUT(10,1)>  …  <OUTPUT${}_{(10,N_{10})}$>
<INPUT11>
To perform text decomposition, <TASK-PROMPT> clarifies the goal to ChatGPT, which involves extracting semantically complete actions from the given text. <INPUT1-10> and <OUTPUT1-10> are filled with 10 pairs of human-crafted examples, each containing several action events extracted from text sampled from Wikitext and BookCorpus. ChatGPT is expected to learn from these examples and use them as a guide to extract action events (<OUTPUT(11,1-N)>) from the final input text (<INPUT11>).
For component extraction, we adjust <TASK-PROMPT> to define the task of extracting the seven components from a given event. We populate <INPUT1-10> and <OUTPUT1-10> with 10 pairs of events and seven comma-separated lists of components extracted from the event, each corresponding to one type of components defined in §3. ChatGPT then extracts seven lists of components for the final given event (<INPUT11>). If any type of component is absent, “None” will be generated instead.
4.2 Component Abstraction and Variation
The next step is designed to implement changes within the event by altering its components, extracted from the previous step, by generating their abstractions or numerical variations. Following [64], we guide ChatGPT by modifying <TASK-PROMPT> with the objective of generating abstract concepts for $s,v,o,se$ and numerical variations for $t,l,n$ within a specified event. For each <INPUT1-10> and <OUTPUT1-10> pair, we populate the input with a specific event and one of its components. The output consists of three human-authored component abstractions or numerical variations that align with the event’s context. Subsequently, ChatGPT is tasked with generating three abstractions or numerical variations for the final pair of the given event and a component within the event (<INPUT11>). Replacing the original components in the event with their generated changes forms changed event candidates for the metaphysical event discrimination task.
Figure 3: An overview of our benchmark curation pipeline with running examples.
4.3 Inference Generation
We then collect inferential states of the modified events by similarly instructing ChatGPT to autonomously generate them. For each altered event, we prompt ChatGPT to separately generate one plausible inference and one metaphysical inference. We first modify <TASK-PROMPT> to generate a state that could potentially be caused by the altered event, and populate <INPUT1-10> with 10 modified events and <OUTPUT1-10> with 10 corresponding plausible inferences authored by human experts. ChatGPT is then requested to generate an additional plausible state inference for the given changed event (<INPUT11>). Next, we adjust <TASK-PROMPT> to generate a metaphysical state that is infrequently caused by the changed event in reality, yet remains contextually relevant. We replace <OUTPUT1-10> with 10 metaphysical inferences and then collect a metaphysical inference from ChatGPT. This, along with the generated plausible inference, forms two candidate data entries for each changed event in the metaphysical inference discrimination task.
4.4 Metaphysical Transition Generation
Given that half of the inferential states generated in the previous step remain metaphysical, we then collect the additional changes necessary to transform these states into plausible real-world inferences. We adjust the <TASK-PROMPT> to describe such required changes and populate <INPUT1-10> with 10 pairs of modified events and their corresponding metaphysical inferences. <OUTPUT1-10> are filled with 10 corresponding human-authored changes in events that can render the inferences plausible. Subsequently, ChatGPT generates the required change for the final pair of the modified event and its metaphysical inference (<INPUT11>). Note that the generated change still needs to be one of the seven types we defined in §3. We collect one additional change for each metaphysical inference and use it as a candidate data entry for the last task. However, we discard event and inference pairs that ChatGPT deems impossible to render plausible, even with an additional change.
4.5 Human Annotations
Annotation: Finally, we carry out large-scale human annotations to label candidate data for each task via Amazon Mechanical Turk (AMT). We provide detailed instructions with examples to qualified workers and task them with annotating (1) the plausibility of the changed events generated in §4.2, (2) the plausibility of the plausible/metaphysical inferences produced in §4.3, and (3) the plausibility of the transitions generated in §4.4. We collect five votes for each entry and the majority vote is used as the final label. The overall inter-annotator agreement (IAA) is 81% in terms of pairwise agreement, and the Fleiss kappa [20] is 0.56, indicating sufficient agreement. More details are in Appendix E.
Verification: To verify the quality of our collected labels, we recruit three postgraduate students with rich experience in NLP research to perform a second round annotation. Each of them is asked to annotate a sample of 100 data entries for each task, following the same instructions provided to the AMT annotators. The results are presented in Table 1. On average, 93.67% labels collected from human annotations align with the expert’s vote, demonstrating the reliability of our collected labels.
Table 1: Statistics of the Mars benchmark in comparison against other benchmarks. Meta. refers to three tasks in Mars. Expert. refers to expert verification results.
Dataset / Task #Text #Event #Avg.Token #Train #Dev #Test #Total. #Unlabel. Expert.
AbsATM [23] N/A 7,196 1.060 107,384 12,117 11,503 131,004 372,584 N/A
AbsPyramid [67] N/A 16,944 1.690 176,691 22,050 22,056 220,797 0 N/A
Meta. Event. 9,998 55,190 1.040 96,004 12,013 11,982 119,999 329,540 91.0%
AbsATM [23] N/A 7,196 6.413 65,386 8,403 7,408 81,197 5,921,195 N/A
Meta. Inference. 9,837 35,528 10.40 96,009 12,010 11,981 120,000 497,590 96.5%
Propara [16] 9,051 9,051 N/A 7,043 913 1,095 9,051 0 N/A
TRAC [25] 15,000 15,000 N/A 10,000 2,000 3,000 15,000 0 N/A
PlanBench [62] 26,250 26,250 N/A 0 0 26,250 26,250 0 N/A
Meta. Transition. 9,677 31,447 1.810 92,495 11,563 11,560 115,618 273,474 93.5%
Table 2: Number of unique components by type in annotated splits of Mars. #Avg. refers to the average number of unique identified/modified component per event.
Component Type Identified Modified
ME. MI. MT. #Avg. ME. MI. MT. #Avg.
Subject 4,376 3,907 3,507 1.116 3,106 2,950 2,591 1.094
Verb 9,874 8,856 8,061 3.647 4,408 4,146 3,760 3.457
Object 12,645 11,302 9,986 1.760 5,949 5,494 4,865 1.703
Temporal Quantifier 3,003 2,560 2,288 0.472 1,394 1,253 1,110 0.435
Spatial Quantifier 3,866 3,741 3,301 0.459 2,064 1,979 1,718 0.476
Numerical Properties 5,619 4,932 4,355 0.652 3,570 3,353 2,920 0.612
Sub-events 419 385 326 0.040 425 402 332 0.037
Total 39,802 35,683 31,824 8.146 20,916 19,577 17,296 7.814
5 Evaluations and Analysis
5.1 [Uncaptioned image]Mars Statistics
Table 1 presents statistics of the Mars benchmark, which comprises a total of 355,617 annotated data distributed across three tasks. We partition the annotated data into training, development, and testing splits following an 8:1:1 ratio, ensuring there is no overlap of text and events between the different splits to preserve the evaluation’s generalizability. On average, 1.4 tokens are
Figure 4: Hypernym distribution of the top 5,000 popular component variations.
generated to describe changes in action for the metaphysical event and transition discrimination tasks, while 10.4 tokens are used for inferences in the metaphysical inference discrimination task. To the best of our knowledge, we are the first in proposing such a triad of tasks concurrently within a single benchmark. To compare Mars with other datasets, we select those with analogous task objectives for each task and compare them individually (see Appendix F). We find Mars tends to be significantly larger than other benchmarks, covering a broader range of events and providing training sets for evaluating the performance of fine-tuned models.
Table 2 presents detailed statistics on the number of unique identified and modified components by type in the annotated splits of each task. The majority (approximately 80%) of the components focus on the subject, verb, and object, while the remainder (around 20%) concentrate on temporal quantifiers, spatial quantifiers, numerical properties, and sub-events. On average, each annotated event in Mars features 8.15 identified components for changes and 7.81 transitions.
To further illustrate the diverse coverage of events and modified components in Mars, we match each component variation against hypernyms in Probase [72] and plot their distribution according to their number of occurrences in Figure 4. Our results indicate that Mars covers over 170,000 hypernyms in Probase, spanning broad categories such as event, activity, concept, unit, and more.
5.2 Main Evaluations on [Uncaptioned image]Mars
5.2.1 Task Setup and Model Selections
We then experiment with a selection of (L)LMs to investigate their performances on our curated Mars benchmark. Following the task definitions in §3, each task is assessed as a binary classification task. For unbiased evaluations, we use accuracy, ROC-AUC, and Macro-F1 scores as evaluation metrics.
Table 3: Evaluation results (%) of various language models on the testing sets of Mars. The best performances within each method are underlined and the best among all methods are bold-faced.
Methods Backbone Event Inference Transition
Acc AUC Ma-F1 Acc AUC Ma-F1 Acc AUC Ma-F1
Random - 50.00 - 49.56 50.00 - 49.56 50.00 - 49.56
Majority - 60.98 - 37.99 58.56 - 36.93 50.25 - 33.37
PTLM (Zero-shot) RoBERTa-Base 211M 38.60 49.40 27.90 44.30 55.11 30.80 51.13 53.37 38.36
RoBERTa-Large 340M 38.57 50.94 27.83 44.37 56.49 30.73 50.90 53.08 33.92
DeBERTa-Base 214M 60.55 49.41 42.89 50.10 47.57 48.96 49.05 41.32 33.19
DeBERTa-Large 435M 48.27 49.88 45.87 47.73 49.94 44.44 50.73 46.96 46.15
GPT2-XL 1.5B 38.62 51.12 27.93 44.40 51.88 31.45 49.92 48.35 48.09
CAR 435M 54.63 49.34 49.96 48.33 42.85 41.93 52.97 35.05 46.94
CANDLE 435M 51.90 49.12 50.30 46.77 44.03 38.48 53.49 34.95 47.95
VERA 11B 51.82 50.48 48.52 60.97 62.54 59.09 61.31 66.32 61.17
PTLM (Fine-tuned) RoBERTa-Base 211M 63.32 62.76 61.76 69.08 70.54 68.90 71.24 72.73 70.65
RoBERTa-Large 340M 64.22 63.18 62.62 69.04 70.63 68.90 69.68 71.70 68.73
DeBERTa-Base 214M 63.82 63.98 63.39 69.50 70.59 69.31 71.96 73.85 71.17
DeBERTa-Large 435M 64.45 64.16 63.27 69.57 71.15 69.33 72.93 74.00 72.01
GPT2-XL 1.5B 46.68 47.63 46.96 43.70 44.22 30.41 44.57 45.03 45.89
VERA 11B 61.95 61.43 60.81 63.90 66.93 70.84 71.75 74.57 73.27
LLM (Zero-shot) Meta-LLaMa-2-7B 50.64 - 41.41 49.87 - 49.23 50.94 - 50.64
Meta-LLaMa-2-13B 51.50 - 49.48 50.81 - 50.57 50.81 - 50.80
Meta-LLaMa-2-70B 52.40 - 49.03 56.13 - 46.81 48.45 - 48.34
Meta-LLaMa-3-8B 50.62 - 49.12 51.33 - 50.98 51.95 - 51.07
Meta-LLaMa-3-70B 57.41 - 50.59 63.40 - 61.82 60.15 - 60.01
Gemma-1.1-7B 56.88 - 48.53 51.83 - 51.76 49.41 - 45.01
Falcon-7B 54.32 - 49.51 51.77 - 50.30 50.42 - 49.02
Falcon-40B 52.35 - 50.36 49.67 - 49.38 50.27 - 50.22
Mistral-7B 49.90 - 48.94 50.23 - 50.06 51.75 - 51.75
LLM (Fine-tuned) Meta-LLaMa-2-7B 60.10 59.90 59.00 63.51 66.44 62.55 66.06 70.38 65.12
Meta-LLaMa-2-13B 60.67 60.64 60.00 64.61 67.67 63.59 68.22 72.19 66.37
Meta-LLaMa-3-8B 60.06 60.54 59.58 65.76 67.88 65.72 69.83 74.59 68.74
Gemma-1.1-7B 61.23 61.25 60.28 69.24 70.76 69.00 73.30 76.91 69.18
Mistral-7B 60.35 60.77 60.07 66.91 70.06 65.95 71.87 75.47 68.53
LLM (API) ChatGPT 51.00 - 50.35 61.35 - 57.63 60.40 - 60.12
ChatGPT (5-shots) 53.61 - 53.28 58.05 - 57.42 62.40 - 59.35
ChatGPT (COT) 53.20 - 52.61 50.40 - 50.32 49.95 - 49.83
ChatGPT (SC-COT) 53.98 - 53.47 52.47 - 51.99 51.25 - 51.13
GPT4 53.90 - 53.45 51.20 - 50.95 49.41 - 49.33
GPT4 (5-shots) 49.85 - 49.58 51.47 - 51.30 48.88 - 48.73
GPT4 (COT) 51.28 - 50.73 51.49 - 51.35 47.62 - 47.58
GPT4 (SC-COT) 51.97 - 51.26 52.05 - 52.27 48.24 - 48.11
The evaluation of different models are categorized into three types: (1) Zero-shot: We first evaluate several (L)LMs in a zero-shot manner. For small-sized Pre-Trained Language Models (PTLMs), we evaluate RoBERTa [42], DeBERTa-v3 [24], GPT2 [52], CAR [63], CANDLE [64], and VERA [40], following the design of zero-shot question answering [44]. For LLMs, we evaluate LLaMa2, LLaMa3 [60, 61], Gemma [47], Falcon [2], and Mistral [33] using direct zero-shot prompting [51]. (2) Finetuning: We then assess the performance of (L)LMs when fine-tuned on the training set of Mars. For PTLMs, we fine-tune RoBERTa, DeBERTa, GPT2-xl, and VERA. For LLMs, we fine-tune LLaMa2, LLaMa3, Gemma, and Mistral using LoRA [27]. (3) LLM API: Finally, we evaluate the performance of ChatGPT [48] and GPT-4 [49], which represent proprietary LLMs, under zero-shot, five-shots, Chain-of-Thought prompting (COT; [68]), and Self-Consistent COT (SC-COT; [66]) settings. Please find implementation details in Appendix D, multi-task fine-tuning experiments in Appendix G, and few-shot fine-tuning experiments in Appendix H.
5.2.2 Results and Analysis
Evaluation results are reported in Table 3. From the results, we observe that:
(1) Most models exhibit subpar performance under the zero-shot setting. Among PTLMs, only VERA delivers acceptable results across all three tasks, while the rest significantly underperform. This indicates the extreme difficulty of our proposed metaphysical reasoning tasks. Though models fine-tuned on commonsense knowledge and conceptualizations [23, 65], such as CAR and CANDLE, show some improvement compared to their DeBERTa-v3-Large backbone, these performances are still unsatisfactory, even falling below the level of majority voting. For LLMs, LLaMa-3-70B outperforms all other LLMs on the three tasks, making it the best zero-shot model. Nevertheless, all models perform poorly across all tasks in Mars, emphasizing the difficulty of our tasks.
(2) Fine-tuning only offers limited benefits. With fine-tuning, all models improve significantly. For example, DeBERTa-Large’s accuracy increases by 16.18%, 21.84%, and 22.2% on three tasks, respectively. However, the best results for all tasks are still capped at around 74%, indicating a shared difficulty and significant room for future enhancements. One potential reason for this is that, since we split the data according to the source of text in Wikitext and BookCorpus, the distribution between different splits may differ significantly, as the domain and topics could be diverse from each other.
(3) The GPT series models underperform compared to other LLMs, and COT does not consistently aid performance. Surprisingly, GPT series models fall short when compared to open LLMs, such as LLaMa-3-70B. One possible explanation is that negative examples in Mars are sourced from ChatGPT’s generation and are obtained via post-human annotation. This makes it challenging to discriminate as these negative examples contradict ChatGPT’s internal knowledge. More advanced prompting methods, like COT, tend to negatively impact the models’ performance.
Table 4: Evaluation results (%) of transfering knowledge from CANDLE to aid Mars. The best performances among each method is underlined and best ones among all methods are bold-faced.
Backbone Training Data Event Inference Transition
Acc AUC Ma-F1 Acc AUC Ma-F1 Acc AUC Ma-F1
DeBERTa 435M Zero-shot 58.27 49.88 45.87 47.73 49.94 44.44 50.73 46.96 46.15
CANDLE 57.94 58.22 57.31 59.43 59.03 58.18 62.00 62.19 61.50
Mars 64.45 64.16 63.27 69.57 71.15 69.33 72.93 74.00 72.01
CANDLE + Mars 64.95 64.27 63.74 71.85 73.32 71.64 74.39 77.97 73.30
VERA 11B Zero-shot 41.82 50.48 38.52 60.97 62.54 59.09 61.31 66.32 61.17
CANDLE 57.81 57.24 56.77 56.59 56.08 55.25 59.79 59.88 59.19
Mars 61.95 61.43 60.81 63.90 66.93 70.84 71.75 74.57 73.27
CANDLE + Mars 62.21 61.77 61.17 71.45 74.46 67.61 73.95 77.35 78.26
LLaMa-3 8B Zero-shot 50.62 - 49.12 51.33 - 50.98 51.95 - 51.07
CANDLE 56.47 56.75 56.07 58.29 57.81 57.00 58.74 58.81 58.19
Mars 60.06 60.54 59.58 65.76 67.88 65.72 69.83 74.59 68.74
CANDLE + Mars 60.93 60.80 60.12 69.13 70.84 72.12 74.09 79.38 71.42
5.3 Analysis
5.3.1 Improving Metaphysical Reasoning via Transferring from Conceptualization Taxonomy
Improving the performance of LLMs on Mars requires extensive fine-tuning on large-scale human-annotated data, making it non-trivial. Given that approximately 80% of action changes are executed by modifying a component along with its abstracted concepts, we first study whether exposing LLMs to more conceptualizations and abstract knowledge can enhance their metaphysical reasoning capabilities. For this purpose, we select CANDLE [64] as the knowledge source, which is an automatically constructed knowledge base containing 382K conceptualizations of events and abstract inferential knowledge. We first convert event-conceptualization pairs into the task format of metaphysical event discrimination and reformat commonsense inferential knowledge to align with the objectives of the metaphysical inference and transition discrimination tasks. More details are in Appendix D.2.
Three backbone models are then fine-tuned separately on CANDLE and Mars. Another group is pre-trained on CANDLE before being fine-tuned on Mars. All models are then evaluated on the testing set of Mars, with the results reported in Table 4. From the results, a significant improvement is observed across all tasks when the models are sequentially fine-tuned on CANDLE and Mars, compared to solely fine-tuning on CANDLE or Mars. These findings indicate that the transfer of conceptualizations and abstract knowledge from CANDLE effectively enhances the performance of LMs in metaphysical reasoning tasks. Since CANDLE is constructed by distilling from ChatGPT without human labor, this opens up a scalable and cost-efficient approach to improving the metaphysical reasoning capabilities of LLMs.
5.3.2 Impact of Component Type on LLM’s Performance
We then analyze the performance of LLMs on each component type to understand the reasons for their subpar performance. We select LLaMa-3-8B as the representative model and compare its accuracy on each component type when fine-tuned on Mars and CANDLE + Mars. The results are illustrated in Figure 5. We observe that while pre-training the model on CANDLE consistently enhances performance, LLaMa3 still struggles when reasoning with changes in spatial quantifiers, temporal quantifiers, and numerical properties. This is in line with recent studies that demonstrate weaknesses in temporal and numerical reasoning for LLMs [57, 55]. Another possible reason is that since CANDLE only contains conceptualizations for subjects, verbs, objects, and sub-events in social events, pre-training models on it cannot provide benefits for the aforementioned aspects of change. Moreover, we only observe limited improvement for the metaphysical event discrimination task. This sheds light on future works that could focus on how to further enhance LLM’s metaphysical reasoning capabilities in these weaker dimensions.
Figure 5: Performances by component types of fine-tuned LLaMa3-8B on three tasks of Mars.
5.3.3 Error Analysis of GPT-Series Models
Finally, we select GPT4 as a representative model and conduct a manual analysis to identify the causes of errors by categorizing the mistakes found in their COT responses. We sample 150 COT responses from each task, all of which result in inconsistent results compared to human annotated labels and present our classifications of these errors as follows:
  • Hallucinations: 41.7% of errors are caused by factual or metaphysical hallucinations by GPT4, where it creates a context that accommodates changes in actions and inferences that are not mentioned in the original text [39, 73]. For instance, in the event “The poet enjoys writing poems about western festivals,” GPT4 incorrectly interprets the poet as Du Fu, an ancient Chinese poet. This leads to a conflict when reasoning about his life and the subsequent inference “He was famous in the west,” resulting in faulty reasoning.
  • Confusion between Concepts and Hypernyms: 36.3% errors are attributed to GPT4’s tendency to perceive abstract components within changed actions as hypernyms that fulfill the change, without considering all potential entities within the original concept. For instance, in a modified event, “He jumps down from very high altitude and lands peacefully,” GPT4 interprets very high altitude as a diving platform, deeming it plausible. However, this concept could also encompass high buildings, which would not be suitable for the event.
  • Internal Conflict: 17.7% errors are attributed to internal conflicts within GPT4’s reasoning rationales, as well as inconsistencies between the binary predictions made and the corresponding reasoning rationales.
  • Annotation Error: 4.3% errors are erroneously identified due to incorrect labels, potentially caused by spamming or a misunderstanding of the task by human annotators.
6 Conclusions
In conclusion, this paper proposes Metaphysical Reasoning to delineate the process of reasoning with changes in distribution and construct [Uncaptioned image]Mars as the associated evaluation benchmark in a non-trivial manner. Our extensive experiments show the difficulty of our proposed task, which cannot be easily addressed through advanced prompting and fine-tuning. Further analysis reveals why LMs encounter difficulties with metaphysical reasoning tasks and suggests a potential avenue for improvement. We hope that our work illuminates the path towards achieving conscious processing in LLMs via System II reasoning through effective comprehension of changes in distribution.
Acknowledgments and Disclosure of Funding
The authors of this paper thank Tianqing Fang, Wenxuan Ding, Wei Fan, Jiaxin Bai, and the anonymous reviewers for their constructive feedbacks. The authors of this paper were supported by the NSFC Fund (U20B2053) from the NSFC of China, the RIF (R6020-19 and R6021-20), and the GRF (16211520 and 16205322) from RGC of Hong Kong. We also thank the support from the UGC Research Matching Grants (RMGS20EG01-D, RMGS20CR11, RMGS20CR12, RMGS20EG19, RMGS20EG21, RMGS23CR05, RMGS23EG08).
Appendix A Limitations and Future Works
Though we consider our work to be a fundamental step towards understanding the capabilities of LMs in reasoning with changes in distribution, several limitations still exist that cannot be covered within this single work. Here, we discuss some important limitations that future works can address:
  • Expand current formulation of changes to include more types of components. In our formulation, we model changes within an action by altering a component within that action to describe the change that has occurred. This approach provides a structured ontology for formulating the various changes that can potentially occur within an event. In our work, we focus on seven primary types of changes, covering the subject, verb, object, spatial quantifier, temporal quantifier, numerical properties, and sub-events of the event. While these seven types encompass most of the potential changes, there are other components within an event that can be modified, such as adjectives, adverbs, and prepositional phrases. Despite this, our benchmark curation pipeline is flexible regarding the types of components it can handle. By incorporating human-authored examples, ChatGPT can perform automatic extraction, thereby easily covering a broader range of component types. Consequently, future research can extend our benchmark to include these additional components.
  • Enabling simultaneous modification of multiple components. For simplicity, we consider only one change occurring per event in each data entry. However, it is also possible for multiple changes to occur simultaneously, thereby modifying more than one component. This, however, could lead to a significantly larger dataset, rendering it impractical to construct a benchmark with human annotation.
  • Implement more detailed changes through instantiation. In our work, we use abstractions to model changes in an event with a hierarchical ontology. The higher the level of abstraction, the more changes it can represent. To reflect these abstracted changes back to their original level in the event, future studies could instantiate them to obtain more detailed changes.
  • Reliance of ChatGPT on benchmark curation. Our data construction process relies significantly on ChatGPT, an expensive and proprietary language model used for data collection, as well as human annotation for data verification. Future research could consider utilizing robust open-source language models [53] and general statement plausibility estimators [40] to replace these methods. This could potentially reduce costs and construct larger benchmark.
Appendix B Ethics Statement
B.1 Offensive Content Elimination
Our benchmark curation pipeline, which involves generating content with ChatGPT, necessitates stringent measures to ensure the absence of offensive content in both the prompts and the generated responses. For this purpose, we apply two strategies to eliminate offensive content. First, we use the highest level of Azure AI Content Safety Filter to filter out any content that contains personal privacy, promotes violence, racial discrimination, hate speech, sexual content, or self-harm. If any such unsafe content is detected in the prompts or generated responses, it automatically triggers a system failure, which prevents the inclusion of such data in our dataset. Second, we manually inspect a random sample of 500 data entries from three tasks in [Uncaptioned image]Mars for offensive content. Based on our annotations, we have not detected any offensive content. We thus believe that our dataset is safe.
B.2 Licenses
We will share our code, dataset, and models under the MIT license, thereby granting other researchers free access to our assets for research purposes. If there are any issues with rights violations or anything related to the data license, we will take full responsibility. Other datasets used in this paper, including Wikitext and Bookcorpus, are shared under the CC-SA license, permitting us to use them for research. As for language models, we access all open-source LMs via the Huggingface Hub [70]. All associated licenses permit user access for research purposes, and we have agreed to follow all terms of use.
B.3 Annotations
We conduct large scale human annotations on the Amazon Mechanical Turk (AMT) platform. The annotators are paid on average at an hourly rate of 19 USD, which is comparable to the minimum wages in the US. The selection of these annotators is solely based on their performance on the evaluation set, and we do not collect any personal information about the participants from AMT. For expert verifications, we have secured IRB approval and support from our institution’s department, which allows us to invite expert graduate students to validate the quality of our data. They all agree to participate voluntarily without being compensated. We have made concerted efforts to eliminate offensive content, thereby ensuring that no annotators are offended.
B.4 Societal Impact
Given that [Uncaptioned image]Mars solely contains events, inferences, and transitions that are either metaphysical or plausible, and considering that we have meticulously removed any offensive content, we believe that this paper will not have any negative societal impact.
Appendix C Prompts for Generations and Evaluations with LLMs
C.1 [Uncaptioned image]Mars Benchmark Curation
We first present our prompts used in each step for sequentially instructing ChatGPT to generate candidate data for [Uncaptioned image]Mars.
C.1.1 Text Decomposition and Event Component Extraction
To decompose a lengthy text from the source corpora into several action events, we use the following prompt to instruct ChatGPT.
You are required to decompose the given long sentence into several short yet semantically complete events, each describing an action. An action event refers to those describing an action or a state change that occurs at a specific time and place. The key components of each event should be preserved: including the subject, verb, object, temporal and spatial quantifiers, numerical properties of the subject and objects, and sub-events. Generate one event as a whole sentence per line. You can generate as many events as you need. Below are some examples:
Sentence <i>: In November 2010, after years of planning and development, SpaceX successfully launched their Falcon 9 rocket into orbit for the first time. The launch took place at Cape Canaveral Air Force Station in Florida. The Falcon 9 carried a Dragon spacecraft mock-up, representing a major milestone in SpaceX’s efforts to develop a reliable and cost-effective means of transporting cargo and eventually astronauts to the International Space Station.
Event 1: SpaceX successfully launched their Falcon 9 rocket into orbit for the first time in November 2010.
Event 2: The Falcon 9 carried a Dragon spacecraft mock-up.
Event 3: The launch of the Falcon 9 took place at Cape Canaveral Air Force Station in Florida.
Sentence <N>: In May 1934, following reports of a Japanese spy operating out of Dutch Harbor, the United States Navy dispatched Edwin T. Layton to the Aleutians to investigate the allegations.
We then use the following prompt to extract seven types of components from the decomposed events.
Given a short event, extract these components:
1. Subject: The noun that performs the action in the sentence.
2. Verb: The action word in the sentence.
3. Object: The noun that receives the action of the verb.
4. Temporal Quantifier: The time or time period of the event in the sentence.
5. Spatial Quantifier: The location or spatial extent of the event in the sentence.
6. Numerical Quantities and Properties of Objects: Numerical values describing the number or properties of the subject, object, or sub-events.
7. Sub-events: Complete events that are part of the main event in the sentence.
For each component, if there are more than one, separate them with |. If you cannot find one for a component, generate ‘‘None’’ only. Below are some examples:
Event <i>: After the First Battle of Naktong Bulge, the US Army’s 2nd Infantry Division was moved to defend the Naktong River line.
Subject: US Army’s 2nd Infantry Division
Verb: moved | defend
Object: None
Temporal Quantifier: After the First Battle of Naktong Bulge
Spatial Quantifier: Naktong River line
Quantities and Properties of Objects: None
Sub-events: The US Army’s 2nd Infantry Division was moved | The US Army’s 2nd Infantry Division was moved to defend the Naktong River line.
Event <N>: The University of Colorado created the Department of Medicine in September 1883 in the Old Main building on the Boulder campus.
C.1.2 Component Abstraction and Variation
For each type of component, we customize the prompt according to the nature of the component and whether the changes are implemented via abstraction or numerical variation. Here, we take the subject category with its abstraction as an example.
Given an event and a subject within the event, abstract the given subject in the given sentence into three different concepts. Each concept should be more abstract than the previous one. You are encouraged to be creative, but please ensure the three concepts gradually cover more instances. Below are some examples:
Event <i>: World’s leading scientists announce breakthrough in clean energy technology, revolutionizing global sustainability efforts.
Subject: World’s leading scientists
Concepts: expert, human, organism
Event <N>: A driver is speeding down the highway.
Subject: A driver
C.1.3 Inference Generation
We use different prompts to collect plausible inferential states and metaphysical inferential states for each changed action event. Here, we provide the prompt for generating a metaphysical inference as an example.
Given an action event, generate a short metaphysical if-then inferential statement that describes an inferential state that only occurs in metaphysical space. A state is a condition or situation in which someone or something exists in the past or present that will last for a certain time if no changes occur. An action is a thing that can be done in a time interval that is usually not long. Metaphysical inference is a type of inference that is not based on empirical evidence but rather on the nature of things. It can be a counterfactual inference that is contrary to the facts or reality, meaning that it is usually not true in reality world. Below are some examples:
Event <i>: In 2003, he played a recurring role on two episodes of The Bill.
Metaphysical Inference: Everyone criticizes his performance in the show.
Event <N>: Sam drives down the road with fast speed.
Note that leveraging LLM to perform contextualized abstraction [64, 75] has been shown to result in larger coverage and stronger downstream benefits compared to previous conceptualization methods [23], such as retrieving from a pre-defined concept taxonomy or human annotation. Our method is justifiable and enables large-scale benchmark construction.
C.1.4 Metaphysical Transition Generation
Finally, we use the prompt below to collect the change needed to transition a metaphysical inference into a plausible one.
You will be given an event and its metaphysical inference, meaning that such an inference is impossible or rarely occurring in reality. Please generate a transition that would make the inference plausible or possible in real life. Specifically, you are required to only change a component of the event. The component must be one of the Subject, Verb, Object, Temporal Quantifier, Spatial Quantifier, Numerical Properties of Subject or Objects, and Sub-events of the event. Below are some examples:
Event <i>: The boss of the company is monitoring the employees.
Metaphysical Inference: The boss feels nervous and is expecting a rise.
Transition: employees -> stocks (Object)
Event <N>: The man is being chased by a 100 meters butterfly in the forest.
Metaphysical Inference: The man is not scared and is laughing.
C.2 Main Evaluations on [Uncaptioned image]Mars
To evaluate LLMs on three tasks in [Uncaptioned image]Mars, we show our evaluating prompts in zero-shot scenario in Table 5. These models are consistent across all models’ evaluations for fair comparison. For few-shot evaluations, few shot examples are added after task descriptions and before the prompted test entry. The exemplars are randomly sampled for each different test entry. For COT prompting, we specifically ask LLMs to “think step by step and generate a short rationale to support your reasoning.” Then, we ask it to give an answer based on its generated rationale. The sampling temperature $\tau$ is set to 0.1 by default, and 5 COT responses are sampled with $\tau$ set to 0.7 in the SC-COT setting.
Table 5: Prompts used for evaluating LLMs across three tasks in [Uncaptioned image]Mars in zero-shot scenario. ME. MI., and MT. stand for three tasks, respectively.
Task Prompt
ME.
Given an event, determine whether it is a metaphysical event or not.
A metaphysical event refers to event that is implausible or rarely occurring in reality.
If it is plausible and commonly accepted in the real world, answer yes.
On the contrary, if the event is metaphysical, answer No.
The event you need to discriminate is: <TEST-ENTRY-EVENT>.
Answer Yes or No only with one word:
MI.
Given an assertion that describes a if-then inference, determine whether the inference is plausible or metaphysical.
A plausible inference is an inference that is likely to be true or reasonable based on the information provided in the assertion.
A metaphysical inference is an inference that is not based on empirical evidence but rather on the nature of things,
it rarely occurs in the real world and can be counterfactual or implausible.
The assertion is: <TEST-ENTRY-INFERENCE>.
Answer Yes or No only with one word.
MT.
You are given an event, an inference based on the event that rarely occurs in the real world (a metaphysical inference),
and a transition in the event that would make the inference plausible or possible in the real world,
please determine whether the transition is correct or not in terms of making the inference plausible or possible.
The event is: <TEST-ENTRY-EVENT>.
The inference is: <TEST-ENTRY-INFERENCE>.
The transition is: <TEST-ENTRY-TRANSITION>.
Answer Yes or No only with one word.
Appendix D Implementation Details
This section provides further implementation details for the main evaluations and subsequent analyses.
For all experiments, we use the Huggingface [^1] https://huggingface.co/ Library [70] to build all models. For each LLM, we conduct experiments with both its instruction fine-tuned version (if any) and the original version. The one achieving higher performances will be included in the reported results. For LLaMa2, the model code is meta-llama/Llama-2-7b/13b/70b(-chat)-hf. For LLaMa3, the model code is meta-llama/Meta-Llama-3-8B/70B(-Instruct). For Mistral, we use mistralai/Mistral-7B(-Instruct)-v0.2.
For ChatGPT and GPT4, we access it through Microsoft Azure APIs [^2] https://azure.microsoft.com/en-us/products/ai-services/. The code of the accessed version for ChatGPT is gpt-35-turbo, and for GPT4 is gpt-4. Both models are of the version dated 2024-02-01. The maximum generation length is set to 50 tokens in zero-shot and few-shot settings, while for COT and SC-COT evaluations, the maximum generation length is set at 200 tokens.
All experiments are conducted on eight NVIDIA-V100 (32G) GPUs, with 8E disk space, 48 CPU cores, and 1T memory. Each experiment is repeated three times with different random seeds, and the average performances are reported.
D.1 Main Evaluations on [Uncaptioned image]Mars
First, we add random voting and majority voting as another two baselines for revealing the characteristics of the [Uncaptioned image]Mars benchmark.
To evaluate PTLMs in a zero-shot manner, we adopt the evaluation pipeline used for zero-shot question answering [44, 63]. Specifically, we convert each discrimination data entry into two declarative statements, which serve as natural language assertions corresponding to ‘yes” or “no” options. For instance, when determining whether an event is metaphysical, we generate two assertions: “The event <EVENT> is metaphysical as it’s unlikely to occur in reality,” and “The event <EVENT> is not metaphysical; it’s plausible in reality.” The models are then tasked with computing the loss of each assertion. The assertion with the lowest loss is considered as the model’s prediction. This approach allows any PTLM to be evaluated under classification tasks with an arbitrary number of options or even type classification based on a single assertion. We use the open code library [^3] https://github.com/Mayer123/HyKAS-CSKG as our code base and follow the default hyperparameter settings. For VERA, we follow the exact same implementation [^4] https://github.com/liujch1998/vera [40]. The accessed backbone model is liujch1998/vera, and all other hyperparameter settings follow the default implementation.
For fine-tuning PTLMs, we connect each PTLM backbone with five fully connected classification layers. The entire model is then fine-tuned using a classification objective with cross-entropy loss. We employ a default setting of a learning rate of 5e-6 and a batch size of 64. The models are optimized using an AdamW optimizer [43], with the model’s performance evaluated every 50 steps. We set the maximum sequence lengths for the tokenizers to 70 for all three discriminative subtasks. Early stopping is also implemented to select the best checkpoint when the highest validation accuracy is achieved. To ensure convergence, we train all models with five epochs.
For evaluating LLMs in a zero-shot manner, we transform the input for each task into assertions using natural language prompts, as illustrated in Table 5. The models are then prompted to determine the plausibility of the provided assertions by answering yes or no questions. We parse their responses using pre-defined rules to derive binary predictions. When generating each token, we consider the top 10 tokens with the highest probabilities.
For fine-tuning LLMs, we use LoRA for fine-tuning, and the LoRA rank and $\alpha$ are set to 16 and 32, respectively. We adopt the open code library from LlamaFactory [^5] https://github.com/hiyouga/LLaMA-Factory [78] for model training and evaluation. We similarly use an Adam [36] optimizer with a learning rate of 5e-5 and a batch size of 8. The maximum sequence length for the tokenizer is set at 300. All models are fine-tuned over three epochs, selecting the checkpoint with the highest accuracy on the validation set.
Finally, for evaluating proprietary LLMs, such as ChatGPT and GPT4, we similarly prompt them as with open LLMs. Detailed prompts are explained in Appendix C.2.
D.2 Improving Metaphysical Reasoning via Transferring from Conceptualization Taxonomy
In this section, we elaborate further on how we transform CANDLE into the format of three tasks in [Uncaptioned image]Mars for large-scale pre-training in improving LMs’ metaphysical reasoning abilities.
CANDLE’s data is primarily divided into two sections. The first section comprises conceptualizations of instances or events, which can be reformatted into metaphysical event discrimination. Each data entry in CANDLE represents a conceptualization of an abstracted instance within an event or the abstraction of an entire event. Following our definition in Section 3, we interpret each conceptualization as a change in the event. For each data entry, replacing the original instance with its conceptualization forms a plausible change that could occur in reality. Subsequently, we randomly select negative conceptualizations for an event from conceptualizations of other events that do not share any common words with the anchor event. These negative conceptualizations form metaphysical events. Three models are then pre-trained on four million events, with a balanced ratio of plausible events and metaphysical events. The hyperparameters for fine-tuning all models remain consistent with the implementation details described above in Appendix D.1.
The second part contains the commonsense inferential knowledge of abstracted events, which can be interpreted as inferential states of the modified events. To synchronize with our task structure, we exclusively select relations that imply a state in the inferential knowledge. We obtain negative inference samples in a similar manner by sampling from inference tails of events without common keywords. Subsequently, we pre-train models for both the metaphysical inference discrimination task and the metaphysical transition reasoning task. These models are trained to determine whether the inference is plausible or metaphysical in relation to the altered event. As CANDLE does not include transitions, this approach serves as the most accurate simulation of the metaphysical transition reasoning task. It’s also important to note that CANDLE is exclusively predicated on social events, covering only subject, object, and sub-events as types of abstraction changes. In contrast, [Uncaptioned image]Mars contains a significantly wider array of events, incorporates more types of changes, and also evaluates (L)LMs’ capabilities in discerning what additional change is requisite to instigate a transition. These features make [Uncaptioned image]Mars distinct from tasks in CANDLE.
Figure 6: Our annotation instruction for the workers at the metaphysical inference discrimination task. Workers are provided with both task explanations and detailed examples.
Figure 7: An example of a question that has been released to the worker. Workers are asked to annotate in a four-point Likert scale.
Appendix E Annotation Details
E.1 Worker Selection Protocol
To ensure the high quality of our human annotation, we implement strict quality control measures. Initially, we invite only those workers to participate in our qualification rounds who meet the following criteria: 1) a minimum of 1K HITs approved, and 2) an approval rate of at least 95%. We select workers separately for each task and conduct three qualification rounds per task to identify those with satisfactory performance. In each qualification round, we create a qualification test suite that includes both easy and challenging questions, each with a gold label from the authors. Workers are required to complete a minimum of 20 questions. To qualify, they must achieve an accuracy rate of at least 80% on the qualification test. After our selection process, we chose 36, 24, and 32 workers for three tasks, respectively, from a pool of 481, 377, and 409 unique annotators. On average, our worker selection rate stands at 7.26%. Following the qualification rounds, workers are required to complete another instruction round. This round contains complex questions selected by the authors, and workers are required to briefly explain the answer to each question. The authors will then double-check the explanations provided by the annotators and disqualify those with a poor understanding.
E.2 Annotation Interface
For each task, we provide workers with comprehensive task explanations in layman’s terms to enhance their understanding. We also offer detailed definitions and several examples of each choice to help annotators understand how to make decisions. Each entry requires the worker to annotate using a four-point Likert scale. Workers are asked to rate the plausibility of the given question using such scale, where 1 signifies strong agreement and 4 indicates strong disagreement. We consider annotations with a value of 1 or 2 as plausible and those with a value of 3 or 4 as implausible. A snapshot of our annotation instructions, along with a snapshot showing the question released to the worker, are shown in Figure 6 and Figure 7. To ensure comprehension, we require annotators to confirm that they have thoroughly read the instructions by ticking a checkbox before starting the annotation task. We also manually monitor the performance of the annotators throughout the annotation process and provide feedback based on common errors. Spammers or underperforming workers will be disqualified. The overall inter-annotator agreement (IAA) stands at 81% in terms of pairwise agreement, and the Fleiss kappa [20] is 0.56. These statistics are generally comparable to or slightly higher than those of other high-quality dataset construction works [54, 18, 19, 29], which indicates that the annotators are close to achieving a strong internal agreement.
E.3 Expert Verification
Finally, we enlist the help of three postgraduate students, each with extensive experience in NLP research, to validate the annotations. These students are given the same instructions as those provided to the crowd-sourcing workers and are asked to verify a sample of 100 annotations for each task. The high level of consistency between our expert annotators and the AMT annotators, as demonstrated in Table 1, suggests that our AMT annotation is of high quality.
Appendix F Task-Oriented Benchmark Comparisons
Table 1 shows a comparison of [Uncaptioned image]Mars with several other datasets, underscoring the unique value of [Uncaptioned image]Mars. In this section, we delve deeper into the differences between various benchmarks for each task, and further elaborates on the distinctive characteristics of [Uncaptioned image]Mars.
F.1 Metaphysical Event Discrimination
In the task of metaphysical event discrimination, we compare [Uncaptioned image]Mars with the discriminative event conceptualization task in AbstractATOMIC and the abstraction detection task in AbsPyramid. Both tasks aim to determine whether a concept feasibly represents an instance within an event (instance abstraction) or the entire event (event abstraction). Despite their similarities to the metaphysical event discrimination task, there are several notable differences. Firstly, none of the previous benchmarks encompass instances of temporal quantifiers, spatial quantifiers, and numerical properties in events, thereby limiting their coverage of instances. This sacrifices a large number of potential changes that could occur within events. Secondly, the concepts in their formulation are disorganized, unlike the increasing abstractive granularity collected in [Uncaptioned image]Mars. Lastly, the primary objective of metaphysical event discrimination is to assess a language model’s ability to discern various abstractions as changes in events, rather than merely evaluating their plausibility in representing instances as concepts.
F.2 Metaphysical Inference Discrimination
In the task of metaphysical inference discrimination, [Uncaptioned image]Mars shares a similar objective with the discriminative triple conceptualization task in AbstractATOMIC. Both tasks evaluate the plausibility of the inference of an abstracted event. However, AbstractATOMIC is limited to featuring social events and, consequently, social inferences. It also only contains nine commonsense relations, and all inference tails are sourced from ATOMIC [54], resulting in very limited semantic coverage. Conversely, [Uncaptioned image]Mars covers a broad range of text events and inferences within various contexts, thanks to the robust generative ability of ChatGPT. Mars also features inferences when the same event is conceptualized in different ways by abstracting different components. This unique feature provides additional value in studying the transition of inferences caused by varying abstractions or variations as changes in the event.
F.3 Metaphysical Transition Discrimination
To the best of our knowledge, no previous benchmarks have covered similar task objective as the metaphysical transition discrimination task. The most comparable tasks are those related to reasoning with changes in logical reasoning or planning, which aim to determine the next necessary step to achieve a goal. This is somewhat akin to inferring the required change in an event to make a metaphysical inference plausible. However, previous works primarily rely on game datasets or feature only a limited number of handcrafted examples, which restricts their effectiveness in evaluating a reasoner’s ability to generalize and understand the consequences of changes across broad domains. Mars addresses this limitation by incorporating a variety of events sourced from Wikitext and Bookcorpus. Previous works also tend to focus solely on selecting the next step from a finite set of possible steps, rather than in an open-ended generative manner. Mars, on the other hand, utilizes ChatGPT to gather additional changes that drive transitions, making it significantly more challenging to reason with transitions in an open-world setting. This approach, however, promotes the development of a generalizable agent with System II reasoning capabilities.
Table 6: Evaluation results (%) of LLMs fine-tuned on [Uncaptioned image]Mars under the multi-task setting.
Methods Backbone Event Inference Transition
Acc AUC Ma-F1 Acc AUC Ma-F1 Acc AUC Ma-F1
Random - 50.00 - 49.56 50.00 - 49.56 50.00 - 49.56
Majority - 60.98 - 37.99 58.56 - 36.93 50.25 - 33.37
LLM (Zero-shot) Meta-LLaMa-2-7B 50.64 - 41.41 49.87 - 49.23 50.94 - 50.64
Meta-LLaMa-2-13B 51.50 - 49.48 50.81 - 50.57 50.81 - 50.80
Meta-LLaMa-2-70B 52.40 - 49.03 56.13 - 46.81 48.45 - 48.34
Meta-LLaMa-3-8B 50.62 - 49.12 51.33 - 50.98 51.95 - 51.07
Meta-LLaMa-3-70B 57.41 - 50.59 63.40 - 61.82 60.15 - 60.01
Gemma-1.1-7B 56.88 - 48.53 51.83 - 51.76 49.41 - 45.01
Falcon-7B 54.32 - 49.51 51.77 - 50.30 50.42 - 49.02
Falcon-40B 52.35 - 50.36 49.67 - 49.38 50.27 - 50.22
Mistral-7B 49.90 - 48.94 50.23 - 50.06 51.75 - 51.75
LLM (Fine-tuned) Meta-LLaMa-2-7B 60.10 59.90 59.00 63.51 66.44 62.55 66.06 70.38 65.12
Meta-LLaMa-2-13B 60.67 60.64 60.00 64.61 67.67 63.59 68.22 72.19 66.37
Meta-LLaMa-3-8B 60.06 60.54 59.58 65.76 67.88 65.72 69.83 74.59 68.74
Gemma-1.1-7B 61.23 61.25 60.28 69.24 70.76 69.00 73.30 76.91 69.18
Mistral-7B 60.35 60.77 60.07 66.91 70.06 65.95 71.87 75.47 68.53
LLM (Multi-task) Meta-LLaMa-2-7B 60.70 59.88 59.17 66.15 64.67 64.34 70.40 70.89 70.20
Meta-LLaMa-2-13B 61.36 61.42 60.69 67.07 66.44 65.68 70.44 69.15 68.62
Meta-LLaMa-3-8B 61.38 61.85 61.02 67.20 67.13 66.60 71.64 72.06 71.12
Gemma-1.1-7B 61.54 62.36 61.15 67.71 67.60 66.98 73.12 72.82 71.89
Mistral-7B 61.03 61.16 60.38 67.69 67.20 66.16 72.34 72.52 71.78
Appendix G Multi-task Fine-tuning on [Uncaptioned image]Mars
G.1 Setup
To achieve conscious processing, an ideal language model should be capable of performing three tasks uniformly and sequentially. However, fine-tuning each task separately contradicts this objective, as it results in a model that can only perform one task after one training. Therefore, in this section, we investigate the possibility of enabling a language model to master all tasks simultaneously through multitask fine-tuning. Given that all three tasks are binary classification tasks, we adopt a straightforward approach. The language model is trained using a randomly shuffled combination of training data from all three tasks. This anticipates that the model will learn all tasks collectively. The best checkpoint is chosen based on achieving the highest accuracy on the validation sets of all three tasks. After training, the model performance is evaluated separately on the testing sets of each task. All training details remain consistent with those explained in the Appendix D.1.
G.2 Results and Analysis
The results are presented in Table 6. Upon analyzing these results, we observe that LLMs fine-tuned in a multi-task setting generally outperform those simply fine-tuned on the respective training data for each task. This observation is interesting as it suggests that training the model uniformly across all three tasks can enhance the entire process simultaneously, thereby improving reasoning with changes in distribution. This implies that LLMs can potentially mimic human learning abilities, which are better equipped to reason with changes by collectively understanding the feasibility, consequence, and necessity of such changes. Such a phenomenon indirectly indicates that our task formulation is indeed interconnected and collectively forms a reasoning pipeline. However, it’s important to note that this improvement is only marginal. LLMs still exhibit limited metaphysical reasoning ability, particularly in the metaphysical event discrimination task. More advanced methods are still required to enable LLMs to achieve metaphysical reasoning.
Table 7: Evaluation results (%) of LLMs fine-tuned on [Uncaptioned image]Mars under the few-shot setting. Training data refers to the ratio of sampled training data from the full training sets of [Uncaptioned image]Mars.
Backbone Training Data Event Inference Transition
Acc AUC Ma-F1 Acc AUC Ma-F1 Acc AUC Ma-F1
LLaMa-2 7B 20% 58.03 58.24 57.62 62.43 64.47 60.43 63.11 63.08 62.73
40% 58.81 58.40 57.69 64.03 67.48 61.58 66.44 70.04 64.15
60% 59.09 59.41 58.62 64.75 68.10 62.79 67.00 70.85 64.15
80% 59.48 60.54 59.82 64.15 68.01 61.53 66.42 70.64 64.92
100% 60.10 59.90 59.00 63.51 66.44 62.55 66.06 70.38 65.12
LLaMa-2 13B 20% 59.95 59.75 58.57 63.80 66.86 61.80 64.11 68.73 64.08
40% 59.45 59.18 58.25 65.49 68.98 63.54 68.52 71.61 64.82
60% 60.19 59.46 58.92 65.90 69.59 64.18 68.24 72.17 65.59
80% 60.24 60.05 59.43 65.99 69.70 64.27 68.35 72.43 65.97
100% 60.67 60.64 60.00 64.61 67.67 63.59 68.22 72.19 66.37
LLaMa-3 8B 20% 60.56 59.91 58.99 63.40 66.77 61.06 65.23 70.50 64.60
40% 60.68 59.98 59.23 62.35 69.00 61.81 69.43 72.72 65.27
60% 60.74 60.88 60.49 65.90 69.59 61.81 69.00 72.78 65.55
80% 60.91 61.03 60.29 66.73 69.71 61.72 68.71 73.15 66.43
100% 60.06 60.54 59.58 65.76 67.88 65.72 69.83 74.59 68.74
Gemma-v1.1 7B 20% 59.07 59.54 59.18 64.70 70.42 62.43 68.41 73.64 67.08
40% 60.79 59.93 59.72 62.80 70.57 62.26 69.83 73.91 62.18
60% 59.26 60.31 59.25 67.83 70.22 60.56 70.68 74.56 66.98
80% 59.31 59.32 58.73 64.03 70.77 63.73 69.66 73.51 67.05
100% 61.23 61.25 60.28 69.24 70.76 69.00 73.30 76.91 69.18
Mistral-v1.1 7B 20% 60.67 60.27 59.61 65.28 69.22 63.16 68.37 72.85 66.15
40% 60.53 60.78 60.03 65.92 70.21 63.96 69.79 72.97 69.46
60% 61.82 61.86 61.07 67.65 70.46 64.09 67.92 73.38 66.76
80% 59.35 59.55 58.85 68.07 70.43 66.49 69.84 73.63 65.84
100% 60.35 60.77 60.07 66.91 70.06 65.95 71.87 75.47 68.53
Appendix H Few-shot Fine-tuning on [Uncaptioned image]Mars
H.1 Setup
From the main evaluation results in Table 3, it is evident that fine-tuning consistently enhances the performance of all models on [Uncaptioned image]Mars. In this section, we delve deeper into the impact of fine-tuning in a few-shot setting, with the aim of analyzing the performance of models trained with limited data. More specifically, we aim to examine how models perform with varying sizes of training data. This will enable us to determine whether collecting more data invariably benefits fine-tuning, thereby leading to the development of more robust metaphysical reasoners. To achieve this, we sample the training data for each task in a progressively increasing ratio of 0.2, 0.4, 0.6, 0.8, and 1.0, and use each sampled training data to fine-tune LLMs for each task individually. The models are then evaluated on the complete validation sets to select the optimal checkpoint, and on the full testing set for performance assessment. All fine-tuning parameters remain consistent across all models, as detailed in Appendix D.1.
H.2 Results and Analysis
The results are reported in Table 7. From these results, we observe that training the model with a few-shot training data sample generally has a negative impact across all tasks in [Uncaptioned image]Mars. However, this impact is not significant, and on rare occasions, the sampled training data even leads to superior results compared to training on the full sets. When the training data is reduced to different ratios (80%, 60%, 40%, and 20%), the performance of the models is not significantly affected. This suggests that the models are capable of learning from a small amount of training data, and the performance is not significantly influenced by the size of the training data. In other words, annotating more data for training does not necessarily result in better performance, indicating that our task cannot be simply resolved by increasing the amount of training data. Future research can explore more advanced reasoning paradigms or training methods to further enhance the capabilities of LLMs in metaphysical reasoning.